# From Rankings to Insights: Evaluation Should Shift Focus from Leaderboard to Feedback

### v2

@@ 5 Results on PC2 Pointwise Evaluation
In this section, we conduct the experiments on different LLM-as-a-Judge methods. The experiment details are shown in § B.1.
5.1 Main Results
@@
Method Overall Writing Roleplay Knowledge Mathematics Coding Reasoning
Pairwise Evaluation 0.8176 0.8000 0.8302 0.8716 0.8750 0.8160 0.7672
Vanilla Pointwise Evaluation 0.6116 0.5838 0.7358 0.5229 0.7708 0.6960 0.5259
   + Baseline Answer 0.6525 0.6054 0.6038 0.6881 0.7500 0.6800 0.6466
NCD Pointwise Evaluation (No Pre-Com) 0.6688 0.6378 0.7358 0.6389 0.7917 0.7097 0.6207
PC2 Pointwise Evaluation (Single Model Pre-Com) 0.6792 0.6432 0.7358 0.6697 0.7500 0.6880 0.6810
   + Diverse Model Pre-Com 0.7453 0.7351 0.8491 0.7248 0.7500 0.7440 0.7328
    + Baseline Answer 0.8318 0.7892 0.7925 0.8257 0.8542 0.8720 0.8707
Main Results. As demonstrated in Tab. 2, we can observe the following: (1) Base pairwise evaluation yields superior performance than vanilla pointwise evaluation, which is consistent with previous study [15]; (2) Vanilla pointwise evaluation performs poorly, but introducing a baseline answer significantly improves accuracy. (3) The naive criteria decomposition (NCD) pointwise evaluation outperforms pointwise evaluation, indicating that performing criterion decomposition alone can significantly enhance performance. (4) Introducing a single model for pre-comparison yields only a marginal improvement in performance compared to NCD, indicating that pre-comparison significantly enhances the quality of the extracted metrics. (5) When diverse models are adopted for pre-comparison, there is a significant improvement, especially in open-domain tasks such as writing and roleplay. (6) Further incorporating baseline answer leads to additional significant improvement, even surpassing the pairwise evaluation. Results on more benchmarks are shown in § A.1.
(a) Point scale.
(b) Criteria number.
Figure 5: Ablation study.
Figure 6: Consistency of rankings across 17 models with varying sample sizes.
Ablation Study. As shown in Fig. 6, we observe slight fluctuations in performance across different point scales. Notably, the model demonstrates superior performance when the point scale is smaller. This is intuitive, as multiple evaluations tend to yield more stable scores with a small point scale. As for number of criteria, too few (e.g., 2-3) or too many (e.g., 3-15) criteria can both adversely affect performance. Based on these findings, we use a 1-5 point scale and 3-9 criteria in all experiments.
Determining the Minimum Meaningful Sample Size. If an evaluation dataset has only one sample for the HTML category, can it reliably indicate which models perform well or poorly in this category? Obviously, different samples can lead to inconsistent evaluation results. This raises an important question: What is the minimum number of samples needed per query type for consistent evaluation? To address this, we analyze tags with over 32 samples. For each tag, we randomly select $N$ samples 20 times, rank 17 LLMs based on each sample set, and measure consistency using the average Spearman’s rho coefficient across the 20 rankings. We then average the consistency scores across all tags. As shown in Fig. 6, consistency improves with larger sample sizes. Our Feedbacker-D-V0 dataset contains at least 19 samples per tag, consistently achieving a score above 0.9, which can be considered reliable.
Figure 7: Token cost comparison.
Table 3: Agreement rate on the winner between two similarly performing models with a small data size.
Model Pair Avg. Min. Std. Dev.
qwen-max-2024-10-15 vs. doubao-1-5-pro-32k-250115
ours 81.52% 45.00% 0.1264
win rate 52.88% 15.00% 0.1586
Meta-Llama-3.1-70B-Instruct vs. Qwen2.5-72B-Instruct
ours 89.39% 55.00% 0.1260
win rate 70.76% 5.00% 0.1875
Meta-Llama-3.1-8B-Instruct vs. Qwen2.5-7B-Instruct
ours 81.67% 35.00% 0.1695
win rate 49.24% 20.00% 0.1801
5.2 Comparison with Other Evaluation Methods
Comparison with Pairwise Evaluation. We compare output token cost with pairwise evaluation and our method. From the results in Fig. 7, we observe that as the number of models increases, the token cost burden of pairwise evaluation becomes significant, whereas our method exhibits linear growth. This indicates that our method is more scalable for comparing a larger number of LLMs.
Comparison with Win Rate. To assess the effectiveness of the win rate metric under limited data conditions, we select pairs of models with similar performance and focus on query types over 100 instances. For each query type, we use the evaluation results from all available data as the ground-truth winner. We then perform 20 random samplings, each consisting of 19 instances, and compare the winner obtained from the sampled instances with the ground-truth winner. The average agreement rate is calculated across all query types and samplings. As shown in Tab. 3, the win rate metric performs poorly with small datasets, highlighting its limitations when data is scarce and model differences are subtle. In contrast, our evaluation method provides a more consistent evaluation.
6 Results on Feedbacker
6.1 A Quick Overview of the Evaluation Results.
@@ 6 Results on Feedbacker
To utilize Feedbacker for analysis, we first employ Feedbacker-D-V0 to evaluate 17 mainstream LLMs in this section. The experimental setup is provided in § B.3. An overview of evaluation results is presented in § B.4. For complete results, please refer to the project homepage.
6.2 Exploring Feedbacker: Practical Usage Examples
(a) Rankings of 17 LLMs under Markup Languages.
(b) Fine-grained rankings.
Figure 8: Model rankings for root.coding.Programming Languages.Markup Languages.
In this section, we showcase the usage of Feedbacker to analyze an LLM’s strengths and weaknesses by several examples. These examples demonstrate the effectiveness and potential of Feedbacker.
To begin with, the Automatic Weakness Reporter reveals an interesting anomaly: gpt-4o-2024-11-20 ranked 1st under the fine-grained query tag root.coding.Programming Languages.Markup Languages.XML, while its overall ranking in Tab. 5 is 6th. This suggests an unusual strength in that specific category.
To further explore this phenomenon, we examine the XML’s parent node root.coding.Programming Languages.Markup Languages using the Visualization Tool in the “By Query” view, which displays model rankings across its subdomains. As shown in Fig. 8(a), claude3.7-sonnet-20250219 ranks 1st overall, while gpt-4o-2024-11-20 ranks 5th. This result indicates that gpt-4o-2024-11-20 does not perform as strongly across the entire Markup Languages category—its lead in XML may be an outlier or the result of targeted optimization.
We then adopt the Failure Mode Explorer to get the results in Fig. 8(b), which reveals a notable distinction: claude3.7-sonnet-20250219 maintains consistent high performance across all subdomains (low standard deviation), whereas gpt-4o-2024-11-20’s strong ranking is mainly driven by its exceptional performance in XML. Its scores in other subdomains are relatively low. This indicates that gpt-4o-2024-11-20 may have been specially optimized for XML tasks, but does not exhibit a broad advantage across the entire Markup Languages domain. In contrast, claude3.7-sonnet-20250219 demonstrates stable, and well-rounded performance. This not only exposes a weakness in gpt-4o-2024-11-20, but also suggests a need for broader optimization.
@@ 7 Related Work
Automatic Evaluation Benchmark. We review existing benchmarks here, focusing on their motivations, dataset designs, and evaluation methods. AlpacaEval-2.0 [16] uses 805 curated general-domain data points to assess instruction-following with win rate metric. MT-Bench [15] includes 80 high-quality queries across 8 domains. It adopts pointwise evaluation on a 1–10 scale. Arena-Hard [18] filters 500 challenging Chatbot Arena datasets and uses win rate metric for evaluation. LiveBench [3] manually updates queries in real-time with rule-based scoring to reduce LLM biases. Auto-Arena [19] employs pairwise debates for Elo-based ranking, with a 40-query evaluation set. WildBench [2] updates real user queries to reflect real-world distributions, using both win rate and pointwise evaluation. We discuss the differences between our methods and existing studies in § A.1.
@@ 8 Conclusion
We propose shifting the focus of evaluation from only rankings and leaderboard to providing valueable feedback. To this end, we introduce Feedbacker, an extensible evaluation framework featuring comprehensive and fine-grained queries, along with tools for visualization and analysis. Our proposed PC2 pointwise evaluation achieves superior accuracy compared to pairwise evaluation while retaining pointwise efficiency. This approach effectively balances precision and speed, offering a practical solution for LLM evaluation. We demonstrate that utilizing Feedbacker enables the convenient identification of a model’s strengths and weaknesses, thereby providing valuable feedback.
While our work employs comprehensive, fine-grained queries to deliver feedback, this approach is not the sole option. We believe Feedbacker can serve as a starting point to inspire the community to develop diverse, impactful feedback mechanisms, enhancing the analytical value of evaluations.
We conduct more in-depth discussions including (1) the implications for future research in § A.2, and (2) the limitations of our work in § A.3.
@@ A.1 Distinctions from Prior Works Enabled by Fine-Grained Dataset
Several existing studies, such as SELF-INSTRUCT [42], Nemotron-4-340B [43], WildBench [2], CLEVA [44], and C2LEVA [5], provide fine-grained evaluation of LLM capabilities. However, their taxonomies are relatively coarse and lack extensibility. For instance, C2LEVA classifies LLM abilities into four broad categories—Language, Knowledge, Reasoning, and Harms—and further subdivides Language into only two types: Typo-Fixing and Transliteration. In contrast, our benchmark introduces an unprecedentedly detailed, fine-grained, and extensible taxonomy. Additionally, none of the aforementioned benchmarks provide visualization or analytical tools tailored for fine-grained evaluation, which is a key contribution of our work.
A concurrent study, EvalTree [45], also organizes queries into a tree structure. However, our TaxBuilder differs fundamentally. (1) Unlike EvalTree, which reorganizes existing datasets, our method starts with a fixed, extensible taxonomy and automatically synthesizes data to populate it. (2) TaxBuilder supports multi-dimensional classification based on topic, style, domain, and programming language, whereas EvalTree organizes queries solely by model capability. (3) Our method fully leverages the power of advanced reasoning models, resulting in more accurate classifications. In contrast, EvalTree relies on sentence embeddings—which may lack precision—and K-Means clustering, with cluster labels generated by LLMs. (4) We provide analysis tools to identify model strengths and weaknesses, which are absent in EvalTree. (6) We also addresses sample size requirements for reliable evaluation within subcategories, an important factor not considered by EvalTree.
A.2 Implications for Future Research
In the future, we believe Feedbacker will provide valuable insights to the community.
(1) For model training, Feedbacker can help identify a model’s strengths and weaknesses. By quantifying fine-grained performance, it enables developers to adjust training parameters accordingly, such as data distribution.
(2) For model evaluation, Feedbacker prevents overestimating model capabilities. Its broad coverage across diverse categories allows it to reveal that models excelling in mainstream tasks may still struggle significantly in less common or specialized areas.
(3) For data synthesis, to the best of our knowledge, TaxBuilder offers an unprecedentedly scalable taxonomy that empowers developers to generate comprehensive, diverse and complex synthetic data for model training. For synthetic data, a concurrent work focusing on constructing code feature trees for code generation tasks [46] conveys ideas similar to those of TaxBuilder.
A.3 Limitations of Our Work
Currently, our framework is limited to language models and does not yet support multimodal models [47, 48, 49, 50, 51]. In future work, we plan to introduce a multimodal extension, Feedbacker-Anything. Moreover, our framework currently lacks support for evaluations related to AI mechanisms such as safety, honesty, factuality, hallucination, and fairness [52, 53, 54, 55, 56, 57, 57, 58, 59, 60, 61]. However, our approach is highly extensible and can be easily adapted to include them as new evaluation domains.
@@ B.1 Experimental Setup for Evaluation Method
Models. The P2C pointwise evaluation involves two types of model. In the pre-comparison stage, multiple models are utilized to generate responses. Here we employ gpt-4o-2024-11-20 [28], doubao-1-5-pro-32k-250115 [35], and deepseek-v3-250324 [34] to ensure the diversity of the responses. For judge model, we adopt QwQ-32B [31] due to its lightweight nature and superior reasoning performance. Unless otherwise specified, we use the officially recommended decoding parameters.
@@
Dataset. All our experiments are based on pairwise human preference datasets Feedbacker-HPD and RMB [62]. Pairwise human preference datasets usually contain a prompt and two responses, and the model’s task is to determine which response is better. For pairwise evaluation, the judge model directly selects the better response. For pointwise evaluation, the judge model assigns scores to both responses, with the higher-scoring response being chosen as the winner.
@@
Feedbacker-HPD is a high-quality human preference dataset containing 636 samples. Queries are carefully filtered, and responses are annotated with majority human labels, making the dataset more representative of human preferences. A detailed description of Feedbacker-HPD is provided in § C.4.4. RMB is a recently popular large-scale human preference dataset comprising 10K queries for helpfulness and 7K for harmlessness. We only use the helpfulness subset in paper.
@@
Unless otherwise specified, our experiments are conducted using our own constructed dataset, Feedbacker-HPD, which will be publicly released alongside Feedbacker-D-V0. Results on RMB are presented only in § B.2.
@@
Baseline. We employ pairwise evaluation, vanilla pointwise evaluation, and naive criteria decomposition (NCD) as baselines. NCD refers to generating query-specific criteria based only on the query, without the process of pre-comparison [63, 7, 64]. Additionally, we also tested the impact of various components on P2C pointwise evaluation. "Single Model Pre-Com" refers to using only gpt-4o-2024-11-20 with a temperature of 1, generating three responses. "Diverse Model Pre-Com" involves using gpt-4o-2024-11-20, doubao-1-5-pro-32k-250115, and deepseek-v3-250324 as models for generating auxiliary responses. For the baseline answer, we use gpt-4o-2024-11-20 with a temperature of 1.0 and top-p set to 0.95, generating a single response as the baseline answer. Regardless of whether we use vanilla pointwise evaluation or our method, when adding the baseline answer, we first evaluate it using the method without the baseline answer and include the evaluation result in the prompt. This process helps standardize the evaluation. Detailed instructions can be found in the respective prompts.
@@
Prompt. For PC2 pointwise evaluation, the prompt for pre-comparison-derived criteria extraction is shown in Fig. 9, the prompt for evaluation execution is shown in Fig. 10, the prompt for evaluation execution with baseline answer is shown in Fig. 11. For pairwise evaluation and vanilla pointwise evaluation, we directly use the prompts from MT-Bench [15]. For NCD, the corresponding prompt is displayed in Fig. 12.
@@
PC2 Pointwise Evaluation (Pre-Comparison-derived Criteria Extraction) You are an impartial judge responsible for evaluating the quality of responses provided by different LLMs to a given [question]. Your task is to design a comprehensive evaluation framework that includes clearly defined metrics and their respective weights. You shuold answer step by step. You should answer in English. Please carefully follow these steps:
1. **Analyze Responses**: You must first compare several provided [answers] and identify their differences. The objective of this comparison is to pinpoint distinguishing factors that significantly influence the quality of the responses.
2. **Develop Metrics**: Establish a hierarchical set of evaluation metrics. There should be 3 to 9 primary metrics. Each primary metric should have several detailed sub-metrics to provide specific, measurable criteria for evaluating the responses.
3. **Assign Weights**: Allocate appropriate weights to each metric based on its relative importance in distinguishing the quality of the responses. The weights should be integers, and the sum of all weights should equal 100.
4. **Output Format**: Present the final evaluation framework in a structured list format. You do not need to include the primary metrics; only the secondary metrics are required, in the following format:
<Evaluation_Framework>
1. Description of Secondary Metric 1 | Weight 1
2. Description of Secondary Metric 2 | Weight 2
3. Description of Secondary Metric 3 | Weight 3
<Evaluation_Framework>
[User Question]
{question}
[The Start of Assistant 1’s Answer]
{answer_1}
[The End of Assistant 1’s Answer]
[The Start of Assistant 2’s Answer]
{answer_2}
[The End of Assistant 2’s Answer]
[The Start of Assistant 3’s Answer]
{answer_3}
[The End of Assistant 3’s Answer]
@@
PC2 Pointwise Evaluation (Evaluation Execution) Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider the evaluation system displayed below. Begin your evaluation by providing an explanation for each metric, assessing the response objectively. Score each metric on a scale of 1 to 3, where 1 represents complete failure to meet the criterion and 3 represents perfection. The weights of the metrics must sum to 100, and the final weighted score should be calculated on a scale up to 300, reflecting the weighted sum of the individual scores. After evaluation, you must summarize the results within <The Start of Evaluation Result> and <The End of Evaluation Result>. Below is an example output:
<The Start of Evaluation Result>
Metric 1 | score: [2]
Metric 2 | score: [3]
Metric 3 | score: [1]
Final Weighted Score: [[200]]
<The End of Evaluation Result>
[Question]
{question}
[The Start of Evaluation System]
{eval_system}
[The End of Evaluation System]
[The Start of Assistant’s Answer]
{answer}
[The End of Assistant’s Answer]
@@
PC2 Pointwise Evaluation (Evaluation Execution with Baseline Answer) Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider the evaluation system displayed below. Begin your evaluation by providing an explanation for each metric, assessing the response objectively. Score each metric on a scale of 1 to 3. Use the baseline answer as a baseline; a higher score indicates a better response compared to the baseline answer, while a lower score indicates a worse response. The weights of the metrics must sum to 100, and the final weighted score should be calculated on a scale of 100 to 300, reflecting the weighted sum of the individual scores. After evaluation, you must summarize the results within <The Start of Evaluation Result> and <The End of Evaluation Result>. Below is an example output:
<The Start of Evaluation Result>
Metric 1 | score: [2]
Metric 2 | score: [3]
Metric 3 | score: [1]
Final Weighted Score: [[200]]
<The End of Evaluation Result>
[Question]
{question}
[The Start of Evaluation System]
{eval_system}
[The End of Evaluation System]
[The Start of Baseline Answer]
{answer_baseline}
[The End of Baseline Answer]
[The Start of Evaluation for Baseline Answer]
{critic_baseline}
[The End of Evaluation for Baseline Answer]
[The Start of Assistant’s Answer]
{answer}
[The End of Assistant’s Answer]

### v3

@@ 5 Results on PC2 Pointwise Evaluation
In this section, we demonstrate the main experimental results. The experiment details are shown in § B.1. Results on more benchmark and ablation results are shown in § A.1.
Method / Model Accuracy
Deepseek-R1
   naive 0.5694
   direct metric decomposition 0.6134
   metric decomposition (single model) 0.5974
   metric decomposition (diverse model) 0.6466
   ours 0.6962
Qwen3-32B
   naive 0.5181
   ours 0.6535
claude-3.7-sonnet
   naive 0.5959
   ours 0.7453
gpt-4.1
   naive 0.6116
   ours 0.7201
@@
As shown in Tab. 2, naive pointwise evaluation yields relatively low accuracy across all tested LLMs, confirming its limited reliability as a judging approach. Introducing metric decomposition markedly improves performance, with direct decomposition and single-model pre-comparison offering moderate gains, and diverse-model pre-comparison delivering more substantial improvements. Building upon this, our method further incorporates baseline answers, leading to consistent and significant accuracy gains over all baselines. These results demonstrate that the combination of criteria decomposition, diverse-model pre-comparison, and baseline guidance forms a robust and generalizable evaluation strategy.
6 Results on SCAN
6.1 Overview of Evaluation Results
@@
To utilize SCAN for analysis, we first employ SCAN-D-V0 to evaluate 21 mainstream LLMs in this section. The experimental setup is provided in § B.3. An overview of evaluation results is presented in § B.4. For complete results, please refer to the project homepage.
6.2 Exploring SCAN: GPT-OSS as Examples
Upon the release of a new model, our SCAN framework facilitates a rapid, comprehensive evaluation of its capabilities across diverse domains. Here, we demonstrate SCAN’s utility through an analysis of the recently released GPT-OSS family models.
Overview. The overall evaluation results summarized in Tab. 8 show that GPT-OSS-120B achieves the highest aggregate score, reflecting its strong overall performance. However, it ranks only 7th in roleplay domain, highlighting a notable weakness obscured by the composite metric. Crucially, our framework reveals that GPT-OSS-20B, despite a lower overall rank, performs exceptionally in coding domain (ranking 2nd), which demonstrates its specialized strength. This underscores the effectiveness of our approach in uncovering fine-grained capability profiles and identifying performance that holistic scores alone cannot capture.
Figure 5: Fine-grained performance comparison on coding.Programming Languages.General-purpose Languages.
Figure 6: The top 4 parent nodes exhibiting the largest performance variance across child nodes, automatically identified by our Failure Mode Explorer tool. The symbol ‘#’ denotes rankings.
Fine-grained Coding Performance Analysis. Our SCAN framework enables decomposition of the aggregate coding score by programming language, allowing evaluation of model performance across specific general-purpose languages such as C and Java. The results shown in Fig. 5 reveal substantial performance variation across languages. Notably, GPT-OSS-120B performs strongly in Python, JavaScript, Go, and Rust, but shows weaker performance in C and Java—areas where Deepseek-V3 achieves superior results. This level of granularity is essential for applications demanding language-specific usage. Furthermore, the analysis uncovers the "spiky" capability profile of GPT-OSS-20B: it ranks 2nd in Python and Rust, and outperforms its larger counterpart, GPT-OSS-120B, in both C and C#. This demonstrates that smaller models can excel in specific tasks, providing a more efficient option for domain-specific development. SCAN’s fine granularity thus facilitates informed, use-case-driven model selection, advancing beyond holistic rankings toward nuanced capability profiling.
Diagnosing Knowledge Instability through Fine-Grained Analysis. Here, we demonstrate the use of our analytical tool, Failure Mode Explorer (see § C.3). In Fig. 6, we present the tool’s automatically identified 4 parent nodes with the largest performance variance across sub-nodes for GPT-OSS-120B. Interestingly, all four are within the knowledge domain. Combined with the observation from Tab. 8 that GPT-OSS-120B’s overall performance in knowledge is unexceptional, this motivates a deeper investigation into its capabilities within this domain.
While GPT-OSS-120B ranks third overall in the knowledge domain, a closer look reveals that its performance is far from uniformly strong. In GPT-OSS-120B’s case, the internal performance distribution is highly uneven. For example, within the technical engineering field (parent node ranked #2), the model achieves top positions in computer science and aerospace engineering (#1), yet falls sharply to #11 in bioengineering. This demonstrates that, while aggregate scores facilitate macro-level comparisons, they may conceal important weaknesses in a model’s capabilities.
To tackle these hidden weaknesses, our SCAN framework converts coarse-grained aggregate metrics into fine-grained diagnostic signals that expose exactly where the gaps lie. In the case above, instead of settling for a general verdict like “the model’s knowledge capability is weak,” SCAN can pinpoint the specific sub-domains—such as bioengineering—that are underperforming relative to other strengths. It is worth noting that these results were automatically detected by our Failure Mode Explorer tool, further validating SCAN’s effectiveness in systematically revealing hidden problems.
@@ 7 Related Work
Automatic Evaluation Benchmark. We review existing benchmarks here, focusing on their motivations, dataset designs, and evaluation methods. AlpacaEval-2.0 Dubois et al. (2024) uses 805 curated general-domain data points to assess instruction-following with win rate metric. MT-Bench Zheng et al. (2023) includes 80 high-quality queries across 8 domains. It adopts pointwise evaluation on a 1–10 scale. Arena-Hard Li et al. (2024e) filters 500 challenging Chatbot Arena datasets and uses win rate metric for evaluation. LiveBench White et al. (2024) manually updates queries in real-time with rule-based scoring to reduce LLM biases. Auto-Arena Zhao et al. (2024) employs pairwise debates for Elo-based ranking, with a 40-query evaluation set. WildBench Lin et al. (2024) updates real user queries to reflect real-world distributions, using both win rate and pointwise evaluation. We discuss the differences between our methods and existing studies in § A.1.
@@ 8 Conclusion
In this paper, we propose shifting the focus of evaluation from only rankings and leaderboards to providing valuable insights. To this end, we introduce SCAN, an extensible evaluation framework featuring comprehensive and fine-grained queries, along with tools for visualization and analysis. Additionally, our proposed PC2-based LLM-as-a-Judge achieves superior accuracy compared to classic evaluation methods. Finally, through extensive experiments, we demonstrate that utilizing SCAN enables the convenient identification of a model’s strengths and weaknesses, thereby providing valuable insights.
In the future, we believe SCAN will provide valuable insights to the community in the following aspects. (1) For model evaluation, SCAN helps prevent one-sided understanding of a model’s capabilities, as its broad coverage across diverse categories. (2) For model training, SCAN can help identify a model’s strengths and weaknesses; by quantifying fine-grained performance, it enables developers to adjust training parameters accordingly, such as data distribution. (3) For data synthesis, TaxBuilder offers a scalable taxonomy that empowers developers to generate comprehensive, diverse, and complex synthetic data for model training.
Limitations
Currently, our framework is designed primarily for language models, and does not yet provide support for multimodal models Li et al. (2024d); Patraucean et al. (2023); Li et al. (2024a); Roberts et al. (2025); Liang et al. (2021). In future work, we aim to extend the framework to include multimodal capabilities through a dedicated module, tentatively named SCAN-Anything.
In addition, the current implementation does not encompass evaluation dimensions related to AI mechanisms such as safety, honesty, factuality, hallucination, and fairness Ren et al. (2024); Röttger et al. (2025); Gu et al. (2024b); Jin et al. (2024); Wang et al. (2024); Liu et al. (2025a); Luo et al. (2024); Ravi et al. (2024); Ding et al. (2024); Huang et al. (2023). Nevertheless, the framework is inherently extensible and can be readily adapted to incorporate these aspects as additional evaluation domains in subsequent work.
@@ A.1 Distinctions from Prior Works Enabled by Fine-Grained Dataset
Several existing studies, such as SELF-INSTRUCT Wang et al. (2022), Nemotron-4-340B Adler et al. (2024), WildBench Lin et al. (2024), CLEVA Li et al. (2023), and C2LEVA Li et al. (2024f), provide fine-grained evaluation of LLM capabilities. However, their taxonomies are relatively coarse and lack extensibility. For instance, C2LEVA classifies LLM abilities into four broad categories—Language, Knowledge, Reasoning, and Harms—and further subdivides Language into only two types: Typo-Fixing and Transliteration. In contrast, our benchmark introduces an unprecedentedly detailed, fine-grained, and extensible taxonomy. Additionally, none of the aforementioned benchmarks provide visualization or analytical tools tailored for fine-grained evaluation, which is a key contribution of our work.
A concurrent study, EvalTree Zeng et al. (2025), also organizes queries into a tree structure. However, our TaxBuilder differs fundamentally. (1) Unlike EvalTree, which reorganizes existing datasets, our method starts with a fixed, extensible taxonomy and automatically synthesizes data to populate it. (2) TaxBuilder supports multi-dimensional classification based on topic, style, domain, and programming language, whereas EvalTree organizes queries solely by model capability. (3) Our method fully leverages the power of advanced reasoning models, resulting in more accurate classifications. In contrast, EvalTree relies on sentence embeddings—which may lack precision—and K-Means clustering, with cluster labels generated by LLMs. (4) We provide analysis tools to identify model strengths and weaknesses, which are absent in EvalTree. (6) We also addresses sample size requirements for reliable evaluation within subcategories, an important factor not considered by EvalTree.
@@ B.1 Experimental Setup for Evaluation Method
Models. The PC2 pointwise evaluation involves two types of model. In the pre-comparison stage, multiple models are utilized to generate responses. Here we employ gpt-4o-2024-11-20 OpenAI (2024a), doubao-1-5-pro-32k-250115 ByteDance (2024), and Deepseek-V3-250324 DeepSeek-AI (2024) to ensure the diversity of the responses. For judge model, we adopt Deepseek-R1 DeepSeek-AI (2025) due to its superior reasoning performance. Unless otherwise specified, we use the officially recommended decoding parameters.
@@
Dataset. All our experiments are based on our pairwise human preference datasets SCAN-HPD, RewardBench-v2 Malik et al. (2025) and RMB Zhou et al. (2025). Pairwise human preference datasets usually contain a prompt and two responses, and the model’s task is to determine which response is better. In pointwise evaluation, the judge model assigns scores to both responses, with the higher-scoring response being chosen as the winner.
@@
SCAN-HPD is a high-quality human preference dataset containing 636 samples. Queries are carefully filtered, and responses are annotated with majority human labels, making the dataset more representative of human preferences. A detailed description of SCAN-HPD is provided in § C.4.4. RewardBench-v2 is a widely used benchmark for evaluating judge models, covering several domains: factuality, precise instruction following, math, safety, and focus. It is commonly used to measure and compare the performance of different language models. RMB is another recently popular large-scale human preference dataset comprising 10K queries for helpfulness and 7K for harmlessness. We only use the helpfulness subset in paper.
@@
Unless otherwise specified, our experiments are conducted using our own constructed dataset, SCAN-HPD, which will be publicly released alongside SCAN-D-V0. Results on RewardBench-v2 and RMB are presented only in § B.2.
@@
Baseline. We employ naive pointwise evaluation as main baseline. Additionally, we also tested the impact of various components on P2C pointwise evaluation. Direct metric decomposition refers to generating query-specific criteria based only on the query, without the process of pre-comparison Li et al. (2024g); Qin et al. (2024); Gupta et al. (2024). Metric decomposition (single model) refers to using only gpt-4o with a temperature of 1, generating three responses. Metric decomposition (diverse model) involves using gpt-4o, doubao-1-5-pro, and Deepseek-V3 as models for generating auxiliary responses. Metric decomposition (diverse model) does not adopt the baseline guidance mechanism. For the baseline answer, we use gpt-4o-2024-11-20 with a temperature of 1.0 and top-p set to 0.95, generating a single response as the baseline answer. Regardless of whether we use naive pointwise evaluation or our method, when adding the baseline answer, we first evaluate it using the method without the baseline answer and include the evaluation result in the prompt. This process helps standardize the evaluation. Detailed instructions can be found in the respective prompts.
@@
Prompt. For PC2 pointwise evaluation, the prompt for pre-comparison-derived criteria extraction is shown in Fig. 7, the prompt for evaluation execution is shown in Fig. 8, the prompt for evaluation execution with baseline answer is shown in Fig. 9. For pairwise evaluation and naive pointwise evaluation, we directly use the prompts from MT-Bench Zheng et al. (2023). For direct metric decomposition, the corresponding prompt is displayed in Fig. 10.
@@
PC2 Pointwise Evaluation (Pre-Comparison-derived Criteria Extraction) You are an impartial judge responsible for evaluating the quality of responses provided by different LLMs to a given [question]. Your task is to design a comprehensive evaluation framework that includes clearly defined metrics and their respective weights. You shuold answer step by step. You should answer in English. Please carefully follow these steps: 1. **Analyze Responses**: You must first compare several provided [answers] and identify their differences. The objective of this comparison is to pinpoint distinguishing factors that significantly influence the quality of the responses. 2. **Develop Metrics**: Establish a hierarchical set of evaluation metrics. There should be 3 to 9 primary metrics. Each primary metric should have several detailed sub-metrics to provide specific, measurable criteria for evaluating the responses. 3. **Assign Weights**: Allocate appropriate weights to each metric based on its relative importance in distinguishing the quality of the responses. The weights should be integers, and the sum of all weights should equal 100. 4. **Output Format**: Present the final evaluation framework in a structured list format. You do not need to include the primary metrics; only the secondary metrics are required, in the following format: <Evaluation_Framework> 1. Description of Secondary Metric 1 | Weight 1 2. Description of Secondary Metric 2 | Weight 2 3. Description of Secondary Metric 3 | Weight 3 <Evaluation_Framework> [User Question] {question} [The Start of Assistant 1’s Answer] {answer_1} [The End of Assistant 1’s Answer] [The Start of Assistant 2’s Answer] {answer_2} [The End of Assistant 2’s Answer] [The Start of Assistant 3’s Answer] {answer_3} [The End of Assistant 3’s Answer]
@@
PC2 Pointwise Evaluation (Evaluation Execution) Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider the evaluation system displayed below. Begin your evaluation by providing an explanation for each metric, assessing the response objectively. Score each metric on a scale of 1 to 3, where 1 represents complete failure to meet the criterion and 3 represents perfection. The weights of the metrics must sum to 100, and the final weighted score should be calculated on a scale up to 300, reflecting the weighted sum of the individual scores. After evaluation, you must summarize the results within <The Start of Evaluation Result> and <The End of Evaluation Result>. Below is an example output: <The Start of Evaluation Result> Metric 1 | score: [2] Metric 2 | score: [3] Metric 3 | score: [1] Final Weighted Score: [[200]] <The End of Evaluation Result> [Question] {question} [The Start of Evaluation System] {eval_system} [The End of Evaluation System] [The Start of Assistant’s Answer] {answer} [The End of Assistant’s Answer]
@@
PC2 Pointwise Evaluation (Evaluation Execution with Baseline Answer) Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider the evaluation system displayed below. Begin your evaluation by providing an explanation for each metric, assessing the response objectively. Score each metric on a scale of 1 to 3. Use the baseline answer as a baseline; a higher score indicates a better response compared to the baseline answer, while a lower score indicates a worse response. The weights of the metrics must sum to 100, and the final weighted score should be calculated on a scale of 100 to 300, reflecting the weighted sum of the individual scores. After evaluation, you must summarize the results within <The Start of Evaluation Result> and <The End of Evaluation Result>. Below is an example output: <The Start of Evaluation Result> Metric 1 | score: [2] Metric 2 | score: [3] Metric 3 | score: [1] Final Weighted Score: [[200]] <The End of Evaluation Result> [Question] {question} [The Start of Evaluation System] {eval_system} [The End of Evaluation System] [The Start of Baseline Answer] {answer_baseline} [The End of Baseline Answer] [The Start of Evaluation for Baseline Answer] {critic_baseline} [The End of Evaluation for Baseline Answer] [The Start of Assistant’s Answer] {answer} [The End of Assistant’s Answer]